OpenStudy (anonymous):

find all real zeros of the polynomial function f(x)= -54^4+80x^2

OpenStudy (anonymous):

A. x=0, x=+-4 B. x=0, x=16 C. x=0,x=4 D. x=,x=+-16

OpenStudy (kirbykirby):

are you sure that's the right equation? because none of those are zeros of \(80x^2-54^4\)

OpenStudy (anonymous):

oh sorry the -54 is a -5

OpenStudy (anonymous):

@kirbykirby

OpenStudy (kirbykirby):

none of those are roots still even if that's -5^4

OpenStudy (anonymous):

f(x)= -5x^4+80x^2

OpenStudy (kirbykirby):

oh there's an x.. ok

OpenStudy (kirbykirby):

\(-5x^4+80x^2=-5(x^2)^2+80(x^2)=0\) if you substitute \(x^2=y\) you get a regular quadratic \(-5y^2+80y=0\)

OpenStudy (kirbykirby):

\(5y(-y+16)=0\) This implies \(5y=0~or~-y+16=0\)

OpenStudy (kirbykirby):

Can you solve for y?

OpenStudy (anonymous):

idk how

OpenStudy (anonymous):

1/16?

OpenStudy (kirbykirby):

5y=0 -> y=0 -y+16=0 .... -y=-16....y=16

OpenStudy (kirbykirby):

but we said \(y=x^2\), so we have \(x^2=0\) and \(x^2=16\) then solve for x

OpenStudy (anonymous):

32

OpenStudy (anonymous):

@kirbykirby

OpenStudy (kirbykirby):

\(x^2=16\) means \(x=-4\) or \(x=4\)

OpenStudy (kirbykirby):

since \((-4)^2=16\) and \(4^2=16\)

OpenStudy (anonymous):

So B right/?

OpenStudy (anonymous):

A. x=0, x=+-4 B. x=0, x=16 C. x=0,x=4 D. x=,x=+-16

OpenStudy (kirbykirby):

it would be A